OpenStudy (anonymous):

what is the projection of (3,4) onto (3,1)?

OpenStudy (anonymous):

1.3(3,1)

OpenStudy (amistre64):

a.b/|a|^2 is the scalar to a

OpenStudy (anonymous):

?? its wrong?

OpenStudy (amistre64):

of is b and onto is a a.b = 3.3+4.1 = 9+4 = 13 (3,1)^2 = 9+1 = 10 13/10 (3,1)

OpenStudy (amistre64):

its good

OpenStudy (anonymous):

ok can you just my other answer about the sum?